# GLHF: General Learned Evolutionary Algorithm Via Hyper Functions

### v1

@@ 3.3 Design of GPOM
The overall architecture of GPOM is shown in Fig. 1. The parameters that need to be trained in GPOM are $\theta=\{\theta_{1},\theta_{2}\}$. At time step $t$, the population is $\mathbf{X}^{t}$. Initially, we amalgamate the information from $\mathbf{X}^{t}$ to construct descriptive representations of the population, $\mathbf{H}^{t}$ and $\mathbf{Z}^{t}$. LMM adaptively generates $\mathbf{S}^{t}$ based on $\mathbf{H}^{t}$. The multiplication of $\mathbf{X}^{t}$ and $\mathbf{S}^{t}$ yields $\mathbf{V}^{t}$ (see Eq. (4)). Next, LCM adaptively generates $\mathbf{cr}^{t}$ based on its input $\mathbf{Z}^{t}$, and performs a crossover operation based on $\mathbf{cr}^{t}$ to obtain $\mathbf{U}^{t}$. Finally, SM (Wu et al., 2023), a 1-to-1 selection strategy is executed between $\mathbf{U}^{t}$ and $\mathbf{X}^{t}$ to produce the next-generation population $\mathbf{X}^{t+1}$.
$\begin{split}\mathbf{X}^{t+1}&=SM(\mathbf{X}^{t},\mathbf{U}^{t})\\ &=tile(l_{x>0}(\mathbf{M}_{F^{\prime}}-\mathbf{M}_{F}))\odot\mathbf{X}^{t}\\ &+tile(1-l_{x>0}(\mathbf{M}_{F^{\prime}}-\mathbf{M}_{F}))\odot\mathbf{U}^{t}\end{split}$ (17)
@@
Algorithm 1 MetaGBT
@@
1:  Randomly sample the parameter $\theta$ of GPOM.
@@
3:     Sample $\left|TS\right|$ populations of size $n$ to obtain the population set $pop\leftarrow[\mathbf{X}_{1}^{0},\mathbf{X}_{2}^{0},\cdots,\mathbf{X}_{\left|TS\right|}^{0}]$.
@@
9:           $\mathbf{X}_{i}^{t}\leftarrow GPOM(\mathbf{X}_{i}^{t-1},1|\theta)$.
10:           $loss_{i}^{t}\leftarrow l_{i}(\mathbf{X}^{t}_{i},\mathbf{X}_{i}^{t-1},f_{i},\omega^{i})$.
@@
12:        $\theta\leftarrow$ Update $\theta$ using gradient backpropagation algorithm based on $\frac{1}{\left|TS\right|}\sum\limits_{i}loss_{i}^{t}$.
@@
3.4 Tasks, Loss Function & MetaGBT
GPOM is meticulously crafted as a model amenable to end-to-end training based on gradients. While GPOM necessitates gradient information from the training task during the training phase, it exhibits the ability to tackle BBO problems in the testing phase without relying on any gradient information. To ensure the acquisition of an efficient, highly robust, and broadly generalizable optimization strategy, GPOM undergoes training on a diverse set of tasks. Training on these tasks sequentially poses the risk of domain overfitting, local optima entrapment, and diminished generalization performance. Consequently, we introduce a training methodology named MetaGBT.
Table 1: Training functions. $z_{i}=x_{i}-\omega_{i}$.
ID Functions Range
TF1 $\sum_{i}{|x_{i}-\omega_{i}|}$ $x\in[-10,10],\omega\in[-10,10]$
TF3 $\sum_{i}z_{i}^{2}$ $x\in[-100,100],\omega\in[-50,50]$
TF4 $\max\{|z_{i}|,1\leq i\leq d\}$ $x\in[-100,100],\omega\in[-50,50]$
Tasks
We form a training task set $TS=\{f_{i}({\mathbf{X}|\mathbf{\omega}^{j}})\}$,where $i\in[1,4]$ and $j\in[1,N]$, comprising $4N$ tasks derived from Table 1, where $\mathbf{\omega}_{i}$ denotes the task parameter influencing the function’s landscape offset. Our selection of these functions for the training task is motivated by their diverse landscape features. The specific landscape features encompassed in $TS$ are detailed in Appendix B.
Loss Function
To avoid bias of different output scales in TS, for any function $f_{i}$ in $TS$, we design the normalized loss function $l_{i}(\mathbf{X}^{t},\mathbf{X}^{t-1},f_{i},\omega^{i})$. Equation (18) calculates the average fitness difference between the input and output of GPOM, further normalized within $[0,1]$.
$\mathbf{X}^{t}$ $=GPOM(\mathbf{X}^{t-1},1|\theta)$ (18)
$l_{i}$ $=\frac{\frac{1}{|\mathbf{X}^{t}|}{\sum\limits_{\mathbf{x}\in\mathbf{X}^{t}}}f_{i}(\mathbf{x}|\omega^{i})-\frac{1}{|\mathbf{X}^{t-1}|}{\sum\limits_{\mathbf{x}\in\mathbf{X}^{t-1}}{f_{i}(\mathbf{x}|\omega^{i})}}}{\left|{\frac{1}{|\mathbf{X}^{t-1}|}{\sum\limits_{\mathbf{x}\in\mathbf{X}^{t-1}}f_{i}(\mathbf{x}|\omega^{i})}}\right|}$
MetaGBT
The pseudocode for MetaGBT is presented in Algorithm 1. Initially, we sample the GPOM parameter $\theta$ from a standard normal distribution. The objective of MetaGBT is to iteratively update $\theta$ to bring it closer to the global optimum $\theta^{*}$. In line 2, we sample a population for each task in TS. Lines 3, 4 and 5 involve the resampling of task parameters for all tasks in TS, thereby altering the task landscape, augmenting training complexity, and enhancing the learning of robust optimization strategies by GPOM. The final loss function (line 10) is determined by computing the average of the loss functions for all tasks. Subsequently, in line 12, we update $\theta$ using a gradient-based optimizer, such as Adam (Kingma & Ba, 2014). The trained GPOM is then ready for application in solving an unknown BBO problem, as depicted in Algorithm 2.
Algorithm 2 Driving GPOM to Solve Problem
@@ 3.4 Tasks, Loss Function & MetaGBT
1:  GPOM loads the trained parameter $\theta$.
@@
Figure 2: The critical difference diagram illustrates the performance ranking of seven algorithms across 24 BBOB problems with dimensions $d=30,100$, employing Wilcoxon-Holm analysis (Ismail Fawaz et al., 2019) at a significance level of $p=0.05$. Algorithm positions are indicative of their mean scores across multiple datasets, with higher scores signifying a method consistently outperforming competitors. Thick horizontal lines denote scenarios where there is no statistically significant difference in algorithm performance.
@@ 4.1 Experimental Setup
We test the performance of GPOM on the widely used BBO benchmark and two complex real-world problems (see Appendix C). Selected methods include DE (DE/rand/1/bin) (Das & Suganthan, 2010) and ES (($\mu$,$\lambda$)-ES) as population-based baselines, L-SHADE (Tanabe & Fukunaga, 2014) and CMA-ES (Hansen, 2016) as state-of-the-art population-based BBO methods, and LES (Lange et al., 2023b) and LGA (Lange et al., 2023a) as state-of-the-art POMs. GPOM is trained on $TS$ with $T=100$, $n=100$, and $d=10$. Detailed parameters for all compared methods are provided in Appendix E. Please refer to Appendix D for the reasons for choosing these algorithms.
@@ 4.2 Results
BBOB (Hansen et al., 2021)
We evaluate the generalization ability of GPOM across 24 BBOB functions with dimensions $d=30$ and $d=100$, where optimal solutions are located at $\mathbf{0}$. Figure 2 presents the critical difference diagram comparing all algorithms (refer to Appendix Tables 5 and 7, and Figures 11, 12 and 13 for detailed results). GPOM significantly outperforms all methods, showcasing its efficacy across varying dimensions. Despite being trained solely on TF1-TF4 with $d=10$, GPOM excels in higher dimensions ($d=\{30,100,500\}$), with its performance advantage becoming more pronounced with increasing dimensionality. Particularly on complex problems F21-F24, where global structure is weak, GPOM lags behind LSHADE but surpasses other methods, attributed to its adaptability through fine-tuning. Additional experimental results in Figure 3 (see Appendix F.2 for details) demonstrate GPOM’s superior performance even when optimal solutions are perturbed within the range of $[-1,1]$.
We provide further experimental results in Figure 3 (see Appendix F.2 for more information), where the optimal solutions of the functions have been disturbed in the range of $[-1,1]$. The performance of GPOM still dominates that of other algorithms, as shown in Figure 3.
Bipedal Walker (Brockman et al., 2016)
The Bipedal Walker task involves optimizing a fully connected neural network with $d=874$ parameters over $k=800$ time steps to enhance robot locomotion control. In Fig. 4(a), LSHADE shows ineffectiveness, while CMA-ES, LSHADE, and LGA suffer from premature convergence. Conversely, GPOM achieves stable and swift convergence, ultimately attaining the highest score.
Enduro (Brockman et al., 2016)
Enduro task entails controlling a strategy with $d=4149$ parameters across $k=500$ steps, posing greater difficulty than Bipedal Walker. As depicted in Fig. 4(b), LGA and LES exhibit premature convergence and limited exploration. While CMA-ES initially converges slightly faster than GPOM, the latter maintains a superior balance between exploration and exploitation, outperforming LSHADE.
Figure 5: Results of ablation study. The metric used to evaluate performance is the optimal value of the function found, with smaller values being better. Here, $d=30$.
Figure 6: The impact of training dataset size on the performance of GPOM. $d=100$.
Figure 7: Results of GPOMs with different sizes on BBOB tests ($d=100$).
@@
Figure 8: The Time Cost of GPOM
@@ 4.3 Analysis
The ablation study results for the designed modules are presented in Fig. 5 (refer to Appendix Table 8 for additional details). Configurations include UNTRAINED, representing an untrained GPOM with randomly initialized parameters; NO LMM, where the LMM is excluded, and a simple DE/rand/1/bin mutation operator is employed; NO LCM, indicating the absence of the learnable crossover operation, using only binomial crossover; and NO MASK, signifying the omission of the mask operation described in Eq. (11).
While UNTRAINED yields optimal results for F9 and F16, as an untrained GPOM is inherently an optimization strategy, the adaptability of trained GPOM surpasses the baselines in most scenarios. In simpler tasks with $d=30$, UNTRAINED underperforms, demonstrating the advantage of trained GPOM on more complex tasks. Notably, NO LMM and NO LCM excel on F5, F11, and F19, respectively. This could be attributed to potential overfitting of GPOM to the relatively simple training set. The exclusion of mask operation (NO MASK) significantly diminishes GPOM’s performance, highlighting the importance of the mask for global information sharing and population interaction, crucial for maintaining diversity. All modules contribute to GPOM’s overall performance, with the negative impact on GPOM’s performance ranked as follows: NO MASK $>$ NO LMM $>$ UNTRAINED $>$ NO LCM.
@@
Training Dataset
We further investigate the influence of training data scale on GPOM, employing four additional functions detailed in Appendix Table 4 for experimental analysis, with corresponding results displayed in Fig. 6 (refer to Appendix Table 9 for additional details). Numeric labels, from 1 to 8, signify the inclusion of TF1 to TF8 during GPOM training. The choice of the training set significantly impacts GPOM’s performance. Noteworthy observations emerge: 1) Utilizing TF1-TF4 leads to enhanced algorithm performance as the number of functions increases; 2) Introduction of TF5, TF6, and TF7 provides marginal performance improvement; 3) Incorporating TF8 results in a notable performance decline, suggesting that increased function diversity does not necessarily boost GPOM performance. The complexity of TF8 might induce overfitting, diminishing GPOM’s generalization capability. This underscores the importance of balancing simplicity and complexity when constructing the training set, as overly complex functions can adversely affect performance.
Scale of GPOM
@@
We explore the performance of GPOM at different scales, which is shown in Fig. 7 (refer to Appendix Table 10 for additional details). We increase GPOM’s parameter count by perturbing the hidden layers of each module ($d_{m},d_{c}$). Six models are constructed in ascending order of parameter count, labeled as VS (very small), S (small), M (medium), L (large), VL (very large), and XL (extra large) (details in the Appendix Table 3). XL achieves the best performance, while VS and M also perform well. S exhibits the worst performance, and VL performs worse than L. Two core factors contribute to this phenomenon: the number of parameters and training. We observe a complex relationship between the number of parameters and training difficulty. VS, with the fewest parameters, is the easiest to train and performs well on BBOB. Conversely, XL, with a large number of parameters, exhibits the strongest capability to represent strategies, resulting in the best performance. The performance of XL aligns with our expectations. We obtain the following principles: 1) Larger models can have stronger capabilities but are more challenging to train; 2) Training difficulty and model scale do not exhibit a simple linear relationship, warranting further research; 3) Larger models require more functions for effective training.
@@
We assess the training and test time efficiency of GPOM across various architectures on BBOB ($d=10$) and BBOB ($d=100$) respectively, as illustrated in Figure 8. GPOM demonstrates remarkable efficiency in tackling BBO problems, with negligible training costs relative to its exceptional generalization ability and high performance.

### v2

@@ 3.3 Design of POM
We design LMM and LCM to achieve the generation of sample strategy (that is, generate $\mathbf{S}^{t}$) and crossover strategy (that is, generate $\mathbf{cr}^{t}$), respectively. The overall architecture of POM is shown in Fig. 1. The parameters that need to be trained in POM are $\theta=\{\theta_{1},\theta_{2}\}$. At time step $t$, the population is $\mathbf{X}^{t}$. Initially, we amalgamate the information from $\mathbf{X}^{t}$ to construct descriptive representations of the population, $\mathbf{H}^{t}$ and $\mathbf{Z}^{t}$. LMM adaptively generates $\mathbf{S}^{t}$ based on $\mathbf{H}^{t}$. The multiplication of $\mathbf{X}^{t}$ and $\mathbf{S}^{t}$ yields $\mathbf{V}^{t}$ (see Eq. (3)). Next, LCM adaptively generates $\mathbf{cr}^{t}$ based on its input $\mathbf{Z}^{t}$, and performs a crossover operation based on $\mathbf{cr}^{t}$ to obtain $\mathbf{U}^{t}$. Finally, SM [51], a 1-to-1 selection strategy is executed between $\mathbf{U}^{t}$ and $\mathbf{X}^{t}$ to produce the next-generation population $\mathbf{X}^{t+1}$.
$\mathbf{X}^{t+1}=SM(\mathbf{X}^{t},\mathbf{U}^{t})=tile(l_{x>0}(\mathbf{M}_{F^{\prime}}-\mathbf{M}_{F}))\odot\mathbf{X}^{t}+tile(1-l_{x>0}(\mathbf{M}_{F^{\prime}}-\mathbf{M}_{F}))\odot\mathbf{U}^{t}$ (15)
@@
Algorithm 1 MetaGBT             Algorithm 2 Driving POM to Solve Problem
@@
1:  Randomly sample the parameter $\theta$ of POM.
@@
3:     Sample $\left|TS\right|$ populations of size $n$ to obtain $[\mathbf{X}_{1}^{0},\mathbf{X}_{2}^{0},\cdots,\mathbf{X}_{\left|TS\right|}^{0}]$.
@@
9:           $\mathbf{X}_{i}^{t}\leftarrow POM(\mathbf{X}_{i}^{t-1},1|\theta)$.
10:           $loss_{i}^{t}\leftarrow l_{i}(\mathbf{X}^{t}_{i},\mathbf{X}_{i}^{t-1},f_{i},\omega^{i},\lambda)$.
@@
12:        $\theta\leftarrow$ Update $\theta$ based on $\frac{1}{\left|TS\right|}\sum\limits_{i}loss_{i}^{t}$.
@@
1:  POM loads the trained parameter $\theta$.
@@
3.4 Tasks, Loss Function & MetaGBT
POM is meticulously crafted as a model amenable to end-to-end training based on gradients. While POM necessitates gradient information from the training task during the training phase, it exhibits the ability to tackle BBO problems in the testing phase without relying on any gradient information. To ensure the acquisition of an efficient, highly robust, and broadly generalizable optimization strategy, POM undergoes training on a diverse set of tasks. Training on these tasks sequentially poses the risk of domain overfitting, local optima entrapment, and diminished generalization performance. Consequently, we introduce a training methodology named MetaGBT.
Tasks. We form a training task set $TS=\{f_{i}({\mathbf{X}|\mathbf{\omega}^{j}})\}$,where $i\in[1,5]$ and $j\in[1,N]$, comprising $4N$ tasks derived from Table 3 in appendix, where $\mathbf{\omega}_{i}$ denotes the task parameter influencing the function’s landscape offset. Our selection of these functions for the training task is motivated by their diverse landscape features. The specific landscape features encompassed in $TS$ are detailed in Appendix B.
Loss Function. To avoid bias of different output scales in TS, for any function $f_{i}$ in $TS$, we design the normalized loss function $l_{i}(\mathbf{X}^{t},\mathbf{X}^{t-1},f_{i},\omega^{i},\lambda)$. In Equation (16), $l_{i}^{1}$ calculates the average fitness difference between the input and output of the POM, further normalized within $[0,1]$. This encourages convergence of the algorithm. $l_{i}^{2}$ uses standard deviation to simulate the distribution of the output population, encouraging diversity in the output population. $std(\mathbf{X}^{t},j)$ is the standard deviation of the jth dimension of the population. $\lambda$ is a hyperparameter, and we find that setting it to 0.005 can make model training more stable.
$\mathbf{X}^{t}$ $=POM(\mathbf{X}^{t-1},1|\theta)$ (16)
$l^{1}_{i}$ $=\frac{\frac{1}{|\mathbf{X}^{t}|}{\sum\limits_{\mathbf{x}\in\mathbf{X}^{t}}}f_{i}(\mathbf{x}|\omega^{i})-\frac{1}{|\mathbf{X}^{t-1}|}{\sum\limits_{\mathbf{x}\in\mathbf{X}^{t-1}}{f_{i}(\mathbf{x}|\omega^{i})}}}{\left|{\frac{1}{|\mathbf{X}^{t-1}|}{\sum\limits_{\mathbf{x}\in\mathbf{X}^{t-1}}f_{i}(\mathbf{x}|\omega^{i})}}\right|},\ \ l_{i}^{2}=\frac{\sum\limits_{j=1}^{d}{std(\mathbf{X}^{t},j)}}{d},\ \ l_{i}=l_{i}^{1}-\lambda l_{i}^{2}$
MetaGBT. The pseudocode for MetaGBT is presented in Algorithm 1. Initially, we sample the POM parameter $\theta$ from a standard normal distribution. The objective of MetaGBT is to iteratively update $\theta$ to bring it closer to the global optimum $\theta^{*}$. In line 2, we sample a population for each task in TS. Lines 3, 4 and 5 involve the resampling of task parameters for all tasks in TS, thereby altering the task landscape, augmenting training complexity, and enhancing the learning of robust optimization strategies by POM. The final loss function (line 10) is determined by computing the average of the loss functions for all tasks. Subsequently, in line 12, we update $\theta$ using a gradient-based optimizer, such as Adam [52]. The trained POM is then ready for application in solving an unknown BBO problem, as depicted in Algorithm 1.
Figure 2: The critical difference diagram illustrates the performance ranking of seven algorithms across 24 BBOB problems with dimensions $d=30,100$, employing Wilcoxon-Holm analysis [53] at a significance level of $p=0.05$. Algorithm positions are indicative of their mean scores across multiple datasets, with higher scores signifying a method consistently outperforming competitors. Thick horizontal lines denote scenarios where there is no statistically significant difference in algorithm performance.
@@ 4.1 Experimental Setup
We test the performance of POM on the widely used BBO benchmark and two complex real-world problems (see Appendix C). Selected methods include DE (DE/rand/1/bin) [54] and ES (($\mu$,$\lambda$)-ES) as population-based baselines, L-SHADE [22] and CMA-ES [12] as state-of-the-art population-based BBO methods, and LES [25] and LGA [10] as state-of-the-art POMs. POM is trained on $TS$ with $T=100$, $n=100$, and $d=10$. Detailed parameters for all compared methods are provided in Appendix E. Please refer to Appendix D for the reasons for choosing these algorithms.
@@ 4.2 Results
BBOB [55]. We evaluate the generalization ability of POM across 24 BBOB functions with dimensions $d=30$ and $d=100$, where optimal solutions are located at $\mathbf{0}$. Figure 3 presents the critical difference diagram comparing all algorithms (refer to Appendix Tables 4 and 6, and Figures 11, 12 and 13 for detailed results). POM significantly outperforms all methods, showcasing its efficacy across varying dimensions. Despite being trained solely on TF1-TF4 with $d=10$, POM excels in higher dimensions ($d=\{30,100,500\}$), with its performance advantage becoming more pronounced with increasing dimensionality. Particularly on complex problems F21-F24, where global structure is weak, POM lags behind LSHADE but surpasses other methods, attributed to its adaptability through fine-tuning. TurBO [56] is the Bayesian optimization algorithm with the best performance on BBOB [57]. Under little budget conditions, the performance of POM outperforms that of TurBO in most cases (see Appendix G for details).
Bipedal Walker [58]. The Bipedal Walker task involves optimizing a fully connected neural network with $d=874$ parameters over $k=800$ time steps to enhance robot locomotion control. In Fig. 3(a), LSHADE shows ineffectiveness, while CMA-ES, LSHADE, and LGA suffer from premature convergence. Conversely, POM achieves stable and swift convergence, ultimately attaining the highest score.
Enduro [58]. Enduro task entails controlling a strategy with $d=4149$ parameters across $k=500$ steps, posing greater difficulty than Bipedal Walker. As depicted in Fig. 3(b), LGA and LES exhibit premature convergence and limited exploration. While CMA-ES initially converges slightly faster than POM, the latter maintains a superior balance between exploration and exploitation, outperforming LSHADE.
Figure 4: (a) Results of ablation study. The metric used to evaluate performance is the optimal value of the function found, with smaller values being better. Here, $d=30$. (b) Results of POMs with different sizes on BBOB tests ($d=100$).
Figure 5: The impact of training dataset size on the performance of POM. $d=100$. 1 means that the training set only contains $TF1$, and 2 means that the training set only contains $TF1$ and $TF2$, and so on.
Figure 6: Experimental results of fine-tuning tests. $RFI=\frac{performance\ \ improvement}{performance\ \ of\ \ base\ \ POM}$.
@@
Figure 7: (a) Time cost of POM. (b) Testing cost of baselines and POM.
@@ 4.3 Analysis
The ablation study results for the designed modules are presented in Fig. 4 (a) (refer to Appendix Table 7 for additional details). Configurations include UNTRAINED, representing an untrained POM with randomly initialized parameters; NO LMM, where the LMM is excluded, and a simple DE/rand/1/bin mutation operator is employed; NO LCM, indicating the absence of the learnable crossover operation, using only binomial crossover; and NO MASK, signifying the omission of the mask operation described in Eq. (8).
While UNTRAINED yields optimal results for F9 and F16, as an untrained POM is inherently an optimization strategy, the adaptability of trained POM surpasses the baselines in most scenarios. In simpler tasks with $d=30$, UNTRAINED underperforms, demonstrating the advantage of trained POM on more complex tasks. Notably, NO LMM and NO LCM excel on F5, F11, and F19, respectively. This could be attributed to potential overfitting of POM to the relatively simple training set. The exclusion of mask operation (NO MASK) significantly diminishes POM’s performance, highlighting the importance of the mask for global information sharing and population interaction, crucial for maintaining diversity. All modules contribute to POM’s overall performance, with the negative impact on POM’s performance ranked as follows: NO MASK $>$ NO LMM $>$ UNTRAINED $>$ NO LCM.
@@
Fine-tuning Test
We evaluate the fine-tuned POM’s performance on $TF6$-$TF8$ as detailed in Appendix Table 3. We replace LCM with a standard transformer encoder to obtain more stable experimental results. For $\mathbf{x}_{i}$ and $\mathbf{v}_{i}$, we normalize their features and then concatenate them by dimension to obtain $\mathbf{xv}_{i}\in\mathbb{R}^{d\times 2}$. $\mathbf{xv}_{i}$ and the normalized [fitness, ranking] information of $\mathbf{x}_{i}$ in the parent population are concatenated to obtain $\mathbf{xvf}_{i}\in\mathbb{R}^{(d+1)\times 2}$. Based on this input, the transformer encoder will generate $\mathbf{cv}_{i}$ in Eq. (12). Different numbers of $\omega$ are generated as fine-tuning samples for $TF6$-$TF8$, and Algorithm 1 is used to fine-tune POM for each function. The base POM is initially trained on $TF1$-$TF5$. We calculate the relative performance improvement (RFI) achieved by the fine-tuned POM compared to the base POM, with results displayed in Figure 7. Experimental results indicate that fine-tuning POM leads to significant performance improvements even with a small sample size. The method for obtaining fine-tuning samples is not restricted; for black-box tasks, a surrogate model can be constructed to facilitate fine-tuning.
Size of Training Dataset
Any complex problem can be simulated by a polynomial composed of simple basic function terms. To ensure that the optimization strategy learned by POM has robust generalization ability and performance, we should train POM on a set of basic functions.
First, we tested the impact of increasing the number of basic functions in the training set on model performance. Next, we examined the effect of introducing complex functions into the training set. Functions $TF1-TF5$ are basic simple terms. For example, $TF1$ is an absolute value term, and $TF5$ is a square summation term. Functions $TF6-TF8$ are composite terms composed of several basic functions. For instance, $TF6$ includes both a cumulative multiplier term and a cosine term. The test results are shown in Figure 5 (a) and (b) (see Appendix Table 8 for details), respectively.
Experimental results indicate that increasing the number of basic functions leads to an overall improvement in POM performance, whereas the introduction of composite terms results in a significant performance decline. This aligns with our hypothesis.
Scale of POM
@@
We explore the performance of POM at different scales, which is shown in Fig. 4 (b) (refer to Appendix Table 9 for additional details). We increase POM’s parameter count by perturbing the hidden layers of each module ($d_{m},d_{c}$). Six models are constructed in ascending order of parameter count, labeled as VS (very small), S (small), M (medium), L (large), VL (very large), and XL (extra large) (details in the Appendix Table 2). XL achieves the best performance, while VS and M also perform well. S exhibits the worst performance, and VL performs worse than L. Two core factors contribute to this phenomenon: the number of parameters and training. We observe a complex relationship between the number of parameters and training difficulty. VS, with the fewest parameters, is the easiest to train and performs well on BBOB. Conversely, XL, with a large number of parameters, exhibits the strongest capability to represent strategies, resulting in the best performance. The performance of XL aligns with our expectations. We obtain the following principles: 1) Larger models can have stronger capabilities but are more challenging to train; 2) Training difficulty and model scale do not exhibit a simple linear relationship, warranting further research; 3) Larger models require more functions for effective training.
@@
We assess the training and test time efficiency of POM across various architectures on BBOB ($d=10$) and BBOB ($d=100$) respectively, as illustrated in Figure 7. POM demonstrates remarkable efficiency in tackling BBO problems, with negligible training costs relative to its exceptional generalization ability and high performance.